# The Language of Approval: Identifying the Drivers of Positive Feedback Online

### v1

@@ 7.1. Transforming Linguistic Patterns into Design Levers
The most immediate application is developing real-time guidance tools that nudge users toward empirically-supported linguistic patterns. Tools like Post Guidance (Horta Ribeiro et al., 2025) and ConvoWizard Chang et al. (2022) could be enhanced to guide users beyond merely complying with rules toward contributing in a manner that resonate with communities. This could be done using a large language model in the form of a writing assistance tool, which can analyze and provide real-time feedback to users as they draft their contributions. These interventions should be particularly prominent for newcomers, who face baseline disadvantages ($\approx 6\%$ lower odds of getting a high score) but benefit disproportionately from clarity ($65\%$ vs $40\%$ odds increase for readability). Such nudges must remain optional and easily dismissible to respect user autonomy while providing formative learning opportunities.
@@ 7.2. Early Surfacing as Socio-technical Infrastructure
Our call to HCI researchers and platform developers is to evaluate this capability not as a pure machine-learning achievement but as a potential for socio-technical interventions in community governance. Automated detector models for identification of undesirable content (Chandrasekharan et al., 2019; Zhan et al., 2025; Goyal et al., 2025), as well as systems such as ConvEx (Choi et al., 2023) and Needle (Liu et al., 2025) to aid moderators in finding problematic behavior are increasingly common. Future work should develop similar models and tools for surfacing desirable behavior and study their effectiveness on positively reinforcing users.
@@ 7.4. Advancing Theory, Data, and Methods in Online Community Research
Our work also contributes a comprehensive feature set spanning 100+ linguistic, semantic, and psychosocial attributes that future researchers can directly apply or augment for studying online feedback mechanisms. This feature collection captures multiple facets of textual communication that influence community reception. Researchers can leverage this curated feature set as a foundation for studying feedback dynamics on other platforms, investigating cross-cultural variations, or exploring adjacent phenomena like virality (Chan et al., 2024), engagement (Chan et al., 2026), or community norm formation (Chandrasekharan et al., 2018; Reddy and Chandrasekharan, 2023).
@@ 7.5.1. Temporal Generalizability:
7.5.2. Human Evaluation Gap:
While our models achieve strong predictive performance, we do not validate these predictions against human judgments of post quality. Future research should conduct controlled human evaluation studies where moderators and users rank pairs of posts, comparing these rankings against our model predictions. Such validation would help determine whether our approach captures genuine quality signals or merely reflects voting patterns that may diverge from human perceptions of valuable contributions.
@@ Acknowledgements.
A.G. was supported by compute credits from the OpenAI Researcher Access Program. This work used the Delta system at the National Center for Supercomputing Applications through allocation #240481 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. We thank the members of the Social Computing Lab (SCUBA), especially Jackie Chan, for their feedback and input on this work. We also thank Yiren Liu and Yijun Liu for helpful suggestions.
@@ Appendix A LLM Prompt for Interpreting Principal Components
Prompt for Interpreting Principal Components [SYSTEM PROMPT] You are an expert sociolinguist and data analyst. You will receive 60 Reddit posts that sit at the two extremes of a principal-components (PC) extracted from sentence embeddings. Your task is to propose the three most plausible linguistic, stylistic, or thematic interpretations of this PC.
Each interpretation must: - State a concise PC name ($\leq 5$ words). - Explain the underlying dimension in $\leq 2$ sentences. - Cite at least two short snippets---one from the negative end, one from the positive end---that support the interpretation. If multiple interpretations seem equally valid, rank them by plausibility.
[USER PROMPT] Principal Component Analysis was performed on sentence embeddings of Reddit posts, and for each Principal Component, we selected the 30 lowest-scoring and 30 highest-scoring posts.
Here are the posts: Top-scoring: {{}} Lowest-scoring: {{}}
Propose the TOP THREE interpretations of what linguistic, stylistic, or thematic dimension this PC captures. Output exactly the following JSON structure (no additional keys, no explanation):
      "pc_name": "",
      "description": "",
      "evidence_negative": "",
      "evidence_positive": ""
      "pc_name": "",
      "description": "",
      "evidence_negative": "",
      "evidence_positive": ""
      "pc_name": "",
      "description": "",
      "evidence_negative": "",
      "evidence_positive": ""

### v2

@@ 7.1. Transforming Linguistic Patterns into Design Levers
The most immediate application is developing real-time guidance tools that nudge users toward empirically-supported linguistic patterns. Tools like Post Guidance (Horta Ribeiro et al., 2025) and ConvoWizard (Chang et al., 2022) could be enhanced to guide users beyond merely complying with rules toward contributing in a manner that resonate with communities. This could be done using a large language model in the form of a writing assistance tool, which can analyze and provide real-time feedback to users as they draft their contributions. These interventions should be particularly prominent for newcomers, who face baseline disadvantages ($\approx 6\%$ lower odds of getting a high score) but benefit disproportionately from clarity ($65\%$ vs $40\%$ odds increase for readability). Such nudges must remain optional and easily dismissible to respect user autonomy while providing formative learning opportunities.
@@ 7.2. Early Surfacing as Socio-technical Infrastructure
Our call to HCI researchers and platform developers is to evaluate this capability not as a pure machine-learning achievement but as a potential for socio-technical interventions in community governance. Automated detector models for identification of undesirable content (Chandrasekharan et al., 2019; Zhan et al., 2025; Goyal et al., 2025), as well as systems such as ConvEx (Choi et al., 2023) and Needle (Liu et al., 2026) to aid moderators in finding problematic behavior are increasingly common. Future work should develop similar tools to surface desirable behavior (e.g., within Reddit’s “modqueue” (Bajpai and Chandrasekharan, 2026; Lambert et al., 2025a)) and study their effectiveness on positively reinforcing users.
@@ 7.4. Advancing Theory, Data, and Methods in Online Community Research
Our work also contributes a comprehensive feature set spanning 100+ linguistic, semantic, and psychosocial attributes that future researchers can directly apply or augment for studying online feedback mechanisms. This feature collection captures multiple facets of textual communication that influence community reception. Researchers can leverage this curated feature set as a foundation for studying feedback dynamics on other platforms, investigating cross-cultural variations, or exploring adjacent phenomena like virality (Chan et al., 2024), engagement (Chan et al., 2026), or community norm formation (Chandrasekharan et al., 2018; Reddy and Chandrasekharan, 2023). We hypothesize that while core linguistic features (clarity, toxicity) would likely generalize across platforms, Reddit-specific affordances (threaded discussions, voting) may limit transferability to platforms with different structures.
@@
7.5. Ethical Considerations: Community Rewards !=!= Social Values
Our findings reveal which linguistic patterns drive positive community feedback. However, what communities reward may diverge from what is most valuable, either to individual users or to broader social outcomes (Diakopoulos and Naaman, 2011). We explicitly note that upvotes represent community approval rather than a guarantee of accuracy or quality (Section 3.1), yet communities may systematically reward entertaining over informative content, consensus-reinforcing over challenging perspectives, and discussion-generating over problem-solving contributions. This is not a methodological limitation but rather a fundamental feature of crowd-based feedback systems.
This divergence could create equity concerns. For example, we show question-heavy, help-seeking posts face lower odds of high scores, while newcomers are already disadvantaged. A user genuinely seeking technical help may need to write precisely the style our analysis shows communities penalize. An important consideration here is that optimizing for upvotes !=!= optimizing for utility to the poster. Therefore, we argue that guidance systems should be goal-aware (asking users whether they want visibility, answers, or discussion), optional and dismissible to preserve autonomy, and transparent about trade-offs. Rather than teaching universal reward optimization, platforms should consider algorithmic interventions to surface diverse contribution types, protect help-seeking behavior, and provide newcomer affordances. Our findings are a necessary input to building better systems, but the path from empirical findings to ethical design requires careful consideration of whose interests are served and how to preserve authenticity of participation.
@@ 7.6.1. Temporal Generalizability:
7.6.2. Evaluation Gap:
While our models achieve strong predictive performance, we do not validate these predictions against human judgments of post quality. Future research should conduct controlled human evaluation studies where moderators and users rank pairs of posts, comparing these rankings against our model predictions. Such validation would help determine whether our approach captures genuine quality signals or merely reflects voting patterns that may diverge from human perceptions of valuable contributions. Further, we do not evaluate the potential of transformer-based or large language models on this prediction task (Goyal et al., 2026b).
7.6.3. Sensitivity to Unmeasured Confounding:
Our causal claims are conditional on the assumption that residual linguistic variation is as good as random after controlling for observed covariates. We conducted sensitivity analyses (Appendix C) which suggest that unmeasured confounding would need to be implausibly strong with a partial $R^{2}>0.15$, exceeding our strongest observed confounder at $R^{2}\approx 0.12$) to eliminate key effects, though more fragile findings may be more vulnerable to moderate confounding. Future work could address this through field experiments that randomize linguistic guidance or natural experiments that provide exogenous variation in writing style, but our quasi-experimental approach represents the most rigorous feasible design for studying naturally occurring linguistic choices at scale.
7.6.4. Visual Content:
Our analysis focuses exclusively on textual features and cannot account for visual content (images, videos, memes), which likely plays a substantial role in community feedback, particularly in visually-oriented subreddits where our text-based models underperform.
@@ Acknowledgements.
This research was supported by NSF CAREER IIS-2439433. A.G. was supported by compute credits from the OpenAI Researcher Access Program. Furthermore, this work used the Delta system at the National Center for Supercomputing Applications through allocation #240481 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. We thank the members of the Social Computing Lab (SCUBA), especially Jackie Chan, for their feedback and input on this work. We also thank Yiren Liu and Yijun Liu for helpful suggestions.
@@ Appendix A LLM Prompt for Interpreting Principal Components
Figure 7 shows the LLM prompt used for interpreting the extracted principal components.
@@
Appendix B Temporal Robustness Check
To assess whether our findings are stable across time or reflect period-specific effects, we split our observation window into two non-overlapping periods and re-estimated all models independently:
  • Window 1: May 1 - July 14, 2020 (Baseline: May 1-14; Sampling: May 15 - July 14)
  • Window 2: July 15 - September 30, 2020 (Baseline: July 15-28; Sampling: July 29 - September 30)
For each window, we applied the identical methodology described in Section 3. Specifically, we constructed candidate pools for high score, computed baseline covariates from the 14-day baseline periods, performed risk stratification, and estimated logistic regression models with fixed effects. We then extracted the log-odds ratios for all significant features ($q<0.05$ after BH-FDR correction) from each window and computed the Pearson correlation between the two sets of estimates.
The correlation between log-odds ratios across the two time periods is r=0.764(p<0.001)r=0.764~(p<0.001), indicating substantial consistency in which linguistic features drive positive feedback and the magnitude of their effects. This suggests our findings are not artifacts of a specific moment but reflect more stable patterns in how Reddit communities respond to linguistic variation. While some variation exists likely due to different sample compositions and period-specific events—the core drivers (readability, discussion-generating framing, question-avoidance, toxicity penalties) remain consistent across both time windows.
Appendix C Sensitivity Analysis to Unobserved Confounders
Our causal estimates assume that conditional on observed covariates, residual linguistic variation is as good as random with respect to outcomes. While we built in extensive control for confounders, unmeasured factors could bias results so we assess how strong such confounding would need to be to overturn findings.
We compute partial $R^{2}$ (variance explained after accounting for other variables) for our strongest observed confounders as a baseline:
  • Daily posting rate: $R^{2}=0.08$
  • Daily removal rate: $R^{2}=0.03$
  • Average score on the author’s posts: $R^{2}=0.12$
  • Author’s Reddit account age: $R^{2}=0.06$
These numbers indicate how influential unmeasured variables would need to be to substantially bias results.
Since our setting has 100+ continuous treatments rather than one binary treatment, multiple fixed effects, and risk stratification that standard sensitivity packages don’t accommodate, instead of using a traditional Rosenbaum-styled sensitivity check (Rosenbaum, 2004) we simulate unmeasured confounders of varying strength for each key feature and add them to our model to observe when the estimated effect becomes non-significant. In other words, this analysis approximates how strong an unmeasured confounder, measured by its partial $R^{2}$ with both treatment and outcome would need to be in order to eliminate the observed effect.
We focus on key features of readability, discussion framing, question ratio, and toxicity. We find that in order to reduce effects to null (OR = 1.00) for Readability (OR=1.40), we require $R^{2}\geq 0.15$, for Discussion framing (OR=1.43), we require $R^{2}\geq 0.16$, for Question ratio (OR=0.71), we require $R^{2}\geq 0.13$, and for Toxicity (OR=0.95), we require $R^{2}\geq 0.02.$
Therefore, for these effects (except toxicity), unmeasured confounding would need $R^{2}>0.13$, which is stronger than our most powerful observed confounder with $R^{2}=0.12$. We believe that other plausible confounders such as author mood, external events, topic expertise, etc. are unlikely to be this strong, especially given our extensive controls. However, the toxicity effect is relatively more fragile, requiring only $R^{2}\geq 0.02$ to eliminate effect.